OpenStudy (anonymous):

Hi everyone! Need alot of help on this one... f(x) has the Power Series 5+10x+15x^2+20x^3+25x^4+... where -infinity

OpenStudy (anonymous):

Whoever tackles this one, could you please help me with the steps and the intuition behind them? Thanks!

OpenStudy (anonymous):

I am going to write down the steps and answer, but I need help understanding them...one second...

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

\[\int\limits_{0}^{x} 5+10w+15w^2+20w^3 dw\] then it goes...

jimthompson5910 (jim_thompson5910):

wait...does the series end? or is it infinite?

OpenStudy (anonymous):

\[5w+5w^2+5w^3 | from 0 \to x\]

OpenStudy (anonymous):

\[5x+5x^2+5x^3+5x^4\] this is the final answer

OpenStudy (anonymous):

none of this makes any sense to me! lol

OpenStudy (anonymous):

oh it also gave a hint that said assume the constant of integration is 2...which doesn't help at all

jimthompson5910 (jim_thompson5910):

here's what I think is going on f(x) = 5+10x+15x^2+20x^3+25x^4+... (infinite power series) f(w) = 5+10w+15w^2+20w^3+25w^4+... (infinite power series) f(w) = (5+5n)w^n ... where n ranges from n = 0 to n = infinity int(f(w)dw) = int( (5+5n)w^n ) int(f(w)dw) = (5+5n)/(n+1)w^(n+1) + C int(f(w)dw) = (5(1+n))/(n+1)w^(n+1) + C int(f(w)dw) = 5w^(n+1) + C .. where n goes from n = 0 to n = infinity

jimthompson5910 (jim_thompson5910):

Now evaluate at the endpoints from w = 0 to w = x to get [5x^(n+1) + C] - [5(0)^(n+1) + C] 5x^(n+1) + C - 5(0)^(n+1) - C 5x^(n+1)

jimthompson5910 (jim_thompson5910):

So...